# Bringing Back the Context: Camera Trap Species Identification as Link Prediction on Multimodal Knowledge Graphs

### v3

@@ Abstract
Camera traps are valuable tools in animal ecology for biodiversity monitoring and conservation. However, challenges like poor generalization to deployment at new unseen locations limit their practical application. Images are naturally associated with heterogeneous forms of context possibly in different modalities. In this work, we leverage the structured context associated with the camera trap images to improve out-of-distribution generalization for the task of species identification in camera traps. For example, a photo of a wild animal may be associated with information about where and when it was taken, as well as structured biology knowledge about the animal species. While typically overlooked by existing work, bringing back such context offers several potential benefits for better image understanding, such as addressing data scarcity and enhancing generalization. However, effectively integrating such heterogeneous context into the visual domain is a challenging problem. To address this, we propose a novel framework that reformulates species classification as link prediction in a multimodal knowledge graph (KG). This framework seamlessly integrates various forms of multimodal context for visual recognition. We apply this framework for out-of-distribution species classification on the iWildCam2020-WILDS and Snapshot Mountain Zebra datasets and achieve competitive performance with state-of-the-art approaches. Furthermore, our framework successfully incorporates biological taxonomy for improved generalization and enhances sample efficiency for recognizing under-represented species. [^1] Website: https://osu-nlp-group.github.io/COSMO/
1 Introduction
@@
Human activities are increasingly endangering wildlife species, resulting in a significant global decline in animal populations [37, 19, 2]. Therefore, accurately identifying and tracking wildlife species is vital for preserving ecological biodiversity. Camera traps, digital cameras activated by motion or infrared in natural habitats, have become ecologists’ preferred data collection tool [44, 64, 23]. However, manually sifting through the numerous images they capture is a time-consuming and arduous task for experts. This has led to the increased use of computer vision techniques for species recognition [65, 52, 1, 30, 13]. Yet, a challenge has arisen: many of these models overfit to the backgrounds of their training images, diminishing their effectiveness on images from new locations [7, 39]. This underscores the need for more adaptable species classification models that perform well across diverse contexts.
@@
Building on this, cognitive science research has demonstrated the profound influence of contextual information on human perception and visual recognition processes [5, 45, 4]. Particularly in wildlife monitoring, camera trap images are replete with crucial contextual data, such as where (i.e., camera location coordinates) and when (i.e., timestamps) a photo is taken. Furthermore, the structured knowledge of biology taxonomy (e.g., Open Tree Taxonomy [46]) can also provide valuable context for understanding the species in camera trap images. Such context provides important knowledge that can boost the recognition of visual concepts. For instance, the knowledge that a certain feline image was taken from a camera trap in Africa significantly reduces the likelihood of it representing a tiger. In addition, more robust associations might be learned with the aid of contextual information because the context provides invariable knowledge that is unbiased towards variations in the illuminations or angles of an image. This may help to compensate for domain shifts in species images resulting from such variations and potentially lead to better out-of-distribution (OOD) generalizability [6, 21]. Consequently, the incorporation of contextual information in species identification presents a significant problem worthy of investigation.
@@
Nevertheless, contextual information has been under-exploited in the literature of image classification; standard image classification models [57, 24] often disregard the contextual information tied to images. This is partly due to the heterogeneous nature of the context, which makes it challenging to incorporate contextual information in image classification using a unified learning framework. Contextual information in different modalities (e.g., numerical values, textual descriptions, or structured taxonomies) is usually represented separately from the image in distinct feature spaces. The question of effectively combining features from these different spaces within a unified learning framework remains unanswered. Existing research typically treats all the features as additional input to the classifier via feature vector concatenation [32, 6, 21] or utilizes fusion to obtain aggregate representations [16, 18]. Despite their simplicity, such approaches are incapable of capturing complex structural and semantic relationships between images and various contextual information. Additionally, these approaches assume a uniform availability of contextual information across all images, which is often unrealistic in real-world scenarios. As a result, their flexibility is limited, especially when considering situations where certain images may lack some contextual details, such as coordinates or timestamps, like in camera trap photos.
@@
Towards this end, we propose a new learning framework, COSMO (Classification Of Species using Multimodal cOntext), where we first organize all species images and contextual information as a multimodal knowledge graph (KG) and then reformulate species classification as the standard link prediction task on the KG. Specifically, we consider species images, their corresponding labels (which are available in the training data), and their associated attributes provided in the context as entities within our KG (see Figure 1 for an example). We represent the relationships between these entities as edges in our KG (see a more concrete description of our KG construction in Section 3.2). Our KG is multimodal since its entities are from different modalities. In this context, species classification can be framed as a link prediction task, where the objective is to predict the presence of an edge between an image and its corresponding species label within the KG. This learning framework enables a unified way to incorporate heterogeneous contextual information for species classification. Each form of multimodal information is treated as a type of entity, a first-class citizen of the multimodal KG with its representation computed using a modality-specific encoder. The learning process enables the interaction of different modalities in a joint feature space for robust representation learning. In addition, COSMO demonstrates greater flexibility by not assuming uniform availability of all contextual information, unlike previous methods.
@@
Figure 1: Overview of our framework COSMO. Left: Our multimodal knowledge graph for camera traps and wildlife. Photos from camera traps are jointly represented in the KG with contextual information such as time, location, and structured biology taxonomy. The taxonomy is obtained from Open Tree Taxonomy (OTT) [46] or iNaturalist [25]. Right: In our formulation of species classification as link prediction, the plausibility score $\psi(s,r,o)$ of each (subject, relation, object) triple is computed using a KGE model (e.g., DistMult), where the subject, relation, and object are all first embedded into a vector space. Specifically, for our multimodal KG, we represent visual entities using a ResNet-50 pre-trained on ImageNet and represent numerical entities using an MLP. For categorical entities and relations, we directly represent them with embedding lookups.
@@
We utilize the popular DistMult [67] model as our backbone model for link prediction to instantiate the COSMO framework. To assess the performance of COSMO, particularly in OOD generalization, we conduct experiments on the iWildCam2020-WILDS benchmark [30] and Snapshot Mountain Zebra [48], which are standard datasets for species identification in camera trap photos. They contain naturally occurring wildlife photos associated with metadata. Factors like variation in illumination, camera pose, and motion blur pose challenges for robustness and generalization, making these benchmarks an ideal testbed for assessing our framework’s effectiveness. We further show that COSMO offers a unified framework to incorporate heterogeneous context that further improves the performance.
@@
  • We instantiate this framework for species classification of wildlife images, including the construction of a novel multimodal knowledge graph for this problem that integrates temporospatial information and structured biology knowledge.
@@ 2 Related Work
Species Recognition in Camera Traps. Deep neural networks such as CNNs have been successfully deployed for large-scale recognition of camera trap images [65, 43, 60]. This has paved the way for significant savings in logistics costs for biodiversity conservation. However, training such models often requires enormous amounts of data to perform well. Sadegh Norouzzadeh et al. and Bothmann et al. propose active learning approaches to mitigate the sample inefficiency of training species classification models in such systems. Another challenge arises from the tendency of these models to overfit to the backgrounds present in the training images [7, 39], which limits their deployment to new camera trap locations [60]. Improving robustness to new locations is a significant research challenge [8, 61] leading to the curation of datasets like iWildCam2020-WILDS [30] to test out-of-distribution (OOD) generalization for such systems. Domain adaptation approaches in the literature seek to mitigate this issue by distributionally robust optimization [26, 50] or learning domain invariant features [59]. In contrast, this work helps improve the robustness to new camera trap locations by utilizing a multimodal KG of heterogeneous contexts.
@@
Image Classification with Auxiliary Information. Despite the ubiquity of contextual metadata, the potential of leveraging them for image classification has been largely under-explored. Previous studies have primarily treated metadata as additional input features for classifiers [32, 6, 21], representing a shallow use that fails to capture the intricate relationships between metadata and images. Some works have attempted to model pairwise dependencies between images using heuristics based on metadata, such as shared tags on social media [38, 35] or aggregating information from neighborhood images with similar metadata [28] while disregarding more complex relationships among images, metadata, and labels. Metaformer [18] feeds a sequence of image patches and metadata to a Transformer model for their fusion. Additionally, these methods assume a uniform availability of metadata for all images, which is often not the case in reality due to data scarcity. For instance, the camera trap location coordinates may not be available in some cases due to privacy and security reasons. In our work, we do not assume such uniform availability and build the multimodal KG using available metadata.
@@
Apart from the metadata, external sources of knowledge are also used in image classification. For instance,  Jayathilaka et al. embed each class as a vector based on a hierarchy derived from WordNet [40]. Alsallakh et al. develop a class hierarchy-aware CNN for image classification on ImageNet. Similarly, Bertinetto et al. and Zhang et al. design hierarchy-aware objectives to incorporate taxonomy in image representations. BioCLIP [58] verbalizes the taxonomic hierarchy to train a CLIP-style foundational model for species classification across plants, animals, and fungi. Marino et al. represent images as local subgraphs of Visual Genome [31]. In contrast, COSMO constructs a global KG with both metadata and external knowledge, e.g., taxonomy information from Open Tree Taxonomy, and approaches image classification as link prediction within the KG. Our novel formulation is flexible in handling data scarcity of metadata and enables reasoning over diverse relationships present in the KG.
@@
KG Link Prediction. Most real-world KGs are incomplete. The task of link prediction or knowledge graph completion (KGC) tries to infer missing links given the observed ones. Early approaches for link prediction range from translation-based models [12, 34] and semantic matching models [42, 67] to the ones that leverage neural networks like feedforward neural networks [20], CNNs [17, 41], and Transformer-based models [68, 15, 53]. These methods use a parameterized scoring function based on learned entity and relation embeddings to calculate the plausibility of a particular triple. However, it could be challenging to fully encode the rich semantic information of KGs into such shallow embeddings. To mitigate this, Schlichtkrull et al., Vashishth et al., Yu et al., Pahuja et al. use graph neural networks (GNNs) to encode the rich neighborhood context of entities for link prediction. In our framework, we employ a global multimodal KG, which consists of biological taxonomy and metadata, as the context to enhance OOD generalization.
@@
Multimodal KG Reasoning. Multimodal KGs extend traditional KGs by including entities of different modalities such as categorical data, images, numerical data, etc. KBLRN [22] is a pioneering work in multimodal KG reasoning that uses extra information in the form of relational and numerical features for multimodal KG reasoning. Similarly, IKRL [55] proposes a fusion of linguistic and visual information with structured information for link prediction. MKBE [49] constructs a multimodal KG using numerical, image, and textual information, treating them as entities instead of auxiliary features, for the link prediction task. MR-GCN [66] further extends it by including support for more modalities, e.g., numerical, temporal, textual, visual, and spatial predicate links in the multimodal KG. To provide a more expressive way for interaction between different modalities, IMF [33] uses bilinear pooling to fuse multiple modality features and trains it using contrastive learning on the contextual entity representations. Our work leverages link prediction in a multimodal KG to enable OOD generalization for species classification in camera traps.
3 Methodology
3.1 Preliminaries
@@
Problem Setup. The task entails species recognition for camera trap images amidst distribution shifts. The training and test sets comprise images obtained from disjoint camera traps, enabling the evaluation of out-of-domain generalization. During training, we use the multimodal KG to train our model, while we use just the image to make predictions for inference. The goal is to learn visual representations that are robust to distribution shifts by leveraging the rich structural and semantic information provided by the multimodal KG.
3.2 Building the Multimodal KG
@@
  • Taxonomy: The taxonomy forms the core of the multimodal knowledge graph, connecting distinct species to higher-order taxa. For iWildCam2020-WILDS, we obtain the phylogenetic taxonomy corresponding to the species of interest from Open Tree Taxonomy (OTT)  [46] and manually link it to the species in the dataset. For the Snapshot Mountain Zebra dataset, we utilize the iNaturalist taxonomy [25] mapping provided by www.lila.science.
  • Location: The camera trap images are associated with the GPS coordinates of their source cameras. For iWildCam2020-WILDS, this metadata is available for a portion of the images (67%) and is obfuscated within 1 km. for privacy reasons. Animals demonstrate a preference for particular habitats, and thus, the location context attribute is useful for species recognition.
@@
Figure 1 presents a schematic representation of various contexts in a multimodal KG. For location, time, and taxonomy attributes, the corresponding RDF triples can be represented as (<image>, location, <GPS co-ordinate>), (<image>, time, <timestamp>), and (<taxon_1>, parent, <taxon_2>), respectively.
3.3 Model Architecture
We use DistMult [67], a strong baseline on KGE benchmarks, as our backbone KG embedding model. [^2] Recent work [51] showed that simple baselines like DistMult outperform more sophisticated neural network baselines when trained properly. Note that COSMO is a general framework that can leverage a variety of KG embedding models proposed in the literature. DistMult minimizes a bilinear scoring function between the entity embeddings of subject and object entities. For a given triple $(h,r,t)$, the scoring function of DistMult is defined as
$\psi(h,r,t)=\bm{h}^{T}\bm{W}_{r}\bm{t}=\sum_{i=1}^{d}\bm{h}_{i}\cdot diag(\bm{W}_{r})_{i}\cdot\bm{t}_{i}$ (1)

### v5

# Reviving the Context: Camera Trap Species Classification as Link Prediction on Multimodal Knowledge Graphs

@@ Abstract.
Camera traps are important tools in animal ecology for biodiversity monitoring and conservation. However, their practical application is limited by issues such as poor generalization to new and unseen locations. Images are typically associated with diverse forms of context, which may exist in different modalities. In this work, we exploit the structured context linked to camera trap images to boost out-of-distribution generalization for species classification tasks in camera traps. For instance, a picture of a wild animal could be linked to details about the time and place it was captured, as well as structured biological knowledge about the animal species. While often overlooked by existing studies, incorporating such context offers several potential benefits for better image understanding, such as addressing data scarcity and enhancing generalization. However, effectively incorporating such heterogeneous context into the visual domain is a challenging problem. To address this, we propose a novel framework that transforms species classification as link prediction in a multimodal knowledge graph (KG). This framework enables the seamless integration of diverse multimodal contexts for visual recognition. We apply this framework for out-of-distribution species classification on the iWildCam2020-WILDS and Snapshot Mountain Zebra datasets and achieve competitive performance with state-of-the-art approaches. Furthermore, our framework enhances sample efficiency for recognizing under-represented species. [^1] Our code is available at https://github.com/OSU-NLP-Group/COSMO
Camera Traps, Species Classification, Multimodal Knowledge Graph, KG Link Prediction
conference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Semantic networksccs: Computing methodologies Object recognition
1. Introduction
@@
Human activities are increasingly endangering wildlife species, resulting in a significant global decline in animal populations (Maxwell et al., 2016; Díaz et al., 2019; Almond et al., 2020). Therefore, accurately identifying and tracking wildlife species is vital for preserving ecological biodiversity. Camera traps, digital cameras activated by motion or infrared in natural habitats, have become ecologists’ preferred data collection tool (O’Connell et al., 2011; Wearn and Glover-Kapfer, 2017; Glover-Kapfer et al., 2019). However, manually sifting through the numerous images they capture is a time-consuming and arduous task for experts. This has led to the increased use of computer vision techniques for species recognition (Schneider et al., 2018; Weinstein, 2018; Sadegh Norouzzadeh et al., 2020; Ahumada et al., 2020; Koh et al., 2021; Bothmann et al., 2023). Yet, a challenge has arisen: many of these models overfit to the backgrounds of their training images, diminishing their effectiveness on images from new locations (Beery et al., 2018; Miao et al., 2019; Schneider et al., 2020). This underscores the need for more adaptable species classification models that perform well across diverse contexts.
@@
Building on this, cognitive science research has demonstrated the profound influence of contextual information on human perception and visual recognition processes (Bar, 2004; Oliva and Torralba, 2007; Aminoff et al., 2007). Particularly in wildlife monitoring, camera trap images are replete with crucial contextual data, such as where (i.e., camera location coordinates) and when (i.e., timestamps) a photo is taken. Furthermore, the structured knowledge of biology taxonomy (e.g., Open Tree Taxonomy (OpenTreeofLife et al., 2019)) can also provide valuable context for understanding the species in camera trap images. Such context provides important knowledge that can boost the recognition of visual concepts. For instance, the knowledge that a certain feline image was taken from a camera trap in Africa significantly reduces the likelihood of it representing a tiger. In addition, more robust associations might be learned with the aid of contextual information because the context provides invariable knowledge that is unbiased towards variations in the illuminations or angles of an image. This may help to compensate for domain shifts in species images resulting from such variations and potentially lead to better out-of-distribution (OOD) generalizability (Bargoti and Underwood, 2016; Ellen et al., 2019). Consequently, the incorporation of contextual information in species identification presents a significant problem worthy of investigation.
@@
Nevertheless, contextual information has been under-exploited in the literature of image classification; standard image classification models (Simonyan and Zisserman, 2015; He et al., 2016) often disregard the contextual information tied to images. This is partly due to the heterogeneous nature of the context, which makes it challenging to incorporate contextual information in image classification using a unified learning framework. Contextual information in different modalities (e.g., numerical values, textual descriptions, or structured taxonomies) is usually represented separately from the image in distinct feature spaces. The question of effectively combining features from these different spaces within a unified learning framework remains unanswered. Existing research typically treats all the features as additional input to the classifier via feature vector concatenation (Li et al., 2014; Bargoti and Underwood, 2016; Ellen et al., 2019) or utilizes fusion to obtain aggregate representations (Chu et al., 2019; Diao et al., 2022). Despite their simplicity, such approaches are incapable of capturing complex structural and semantic relationships between images and various contextual information. Additionally, these approaches assume a uniform availability of contextual information across all images, which is often unrealistic in real-world scenarios. As a result, their flexibility is limited, especially when considering situations where certain images may lack some contextual details, such as coordinates or timestamps, like in camera trap photos.
@@
Towards this end, we propose a new learning framework, COSMO (Classification Of Species using Multimodal cOntext), where we first organize all species images and contextual information as a multimodal knowledge graph (KG) and then reformulate species classification as the standard link prediction task on the KG. Specifically, we consider species images, their corresponding labels (which are available in the training data), and their associated attributes provided in the context as entities within our KG (see Figure 1 for an example). We represent the relationships between these entities as edges in our KG (see a more concrete description of our KG construction in Section 3.2). Our KG is multimodal because its entities belong to different modalities. In this context, species classification can be framed as a link prediction task, where the objective is to predict the presence of an edge between an image and its corresponding species label within the KG. This learning framework enables a unified way to incorporate heterogeneous contextual information for species classification. Each form of multimodal information is treated as a type of entity, a first-class citizen of the multimodal KG with its representation computed using a modality-specific encoder. The learning process enables the interaction of different modalities in a joint feature space for robust representation learning. In addition, COSMO demonstrates greater flexibility by not assuming uniform availability of all contextual information, unlike previous methods.
@@
Figure 1. Overview of our framework COSMO. Left: Our multimodal knowledge graph for camera traps and wildlife. Photos from camera traps are jointly represented in the KG with contextual information such as time, location, and structured biology taxonomy. The taxonomy is obtained from Open Tree Taxonomy (OTT) (OpenTreeofLife et al., 2019) or iNaturalist (Horn et al., 2018). Right: In our formulation of species classification as link prediction, the plausibility score $\psi(s,r,o)$ of each (subject, relation, object) triplet is computed using a KGE model (e.g., DistMult), where the subject, relation, and object are all first embedded into a vector space. Specifically, for our multimodal KG, we represent visual entities using a ResNet-50 pre-trained on ImageNet and represent numerical entities using an MLP. For categorical entities and relations, we directly represent them with embedding lookups.
@@
We employ the widely used DistMult (Yang et al., 2015) model as our backbone model for link prediction to instantiate the COSMO framework. To assess the performance of COSMO, particularly in terms of out-of-distribution generalization, we conduct experiments on the iWildCam2020-WILDS benchmark (Koh et al., 2021) and Snapshot Mountain Zebra (Pardo et al., 2021), which are standard datasets for species classification in camera trap photos. They contain naturally occurring wildlife photos associated with metadata. Factors like variation in illumination, camera pose, and motion blur pose challenges for robustness and generalization, making these benchmarks an ideal testbed for assessing our framework’s effectiveness. We show that COSMO offers a unified framework to incorporate heterogeneous context leading to improved species classification performance over existing out-of-distribution generalization approaches.
@@
  • We instantiate this framework for species classification of wildlife images, including the construction of a novel multimodal KG for this problem that integrates spatiotemporal information and structured biology knowledge.
@@ 2. Related Work
Species Recognition in Camera Traps. Deep neural networks such as CNNs have been successfully deployed for large-scale recognition of camera trap images (Weinstein, 2018; Norouzzadeh et al., 2018; Tabak et al., 2019). This has paved the way for significant savings in logistics costs for biodiversity conservation. However, training such models often requires enormous amounts of data to perform well. Sadegh Norouzzadeh et al. and Bothmann et al. propose active learning approaches to mitigate the sample inefficiency of training species classification models in such systems. Another challenge arises from the tendency of these models to overfit to the backgrounds present in the training images (Beery et al., 2018; Miao et al., 2019), which limits their deployment to new camera trap locations (Tabak et al., 2019; Schneider et al., 2020). Improving robustness to new locations is a significant research challenge (Beery et al., 2019; Tabak et al., 2020) leading to the curation of datasets like iWildCam2020-WILDS (Koh et al., 2021) to test OOD generalization for such systems. Domain adaptation approaches in the literature seek to mitigate this issue by distributionally robust optimization (Hu et al., 2018; Qi et al., 2023) or learning domain invariant features (Sun and Saenko, 2016). In contrast, this work helps improve the robustness to new camera trap locations by utilizing a multimodal KG of heterogeneous contexts.
@@
Image Classification with Auxiliary Information. Despite the ubiquity of contextual metadata, the potential of leveraging them for image classification has been largely under-explored. Previous studies have primarily treated metadata as additional input features for classifiers (Li et al., 2014; Bargoti and Underwood, 2016; Ellen et al., 2019), representing a shallow use that fails to capture the intricate relationships between metadata and images. Some works have attempted to model pairwise dependencies between images using heuristics based on metadata, such as shared tags on social media (McAuley and Leskovec, 2012; Long et al., 2019) or aggregating information from neighborhood images with similar metadata (Johnson et al., 2015) while disregarding more complex relationships among images, metadata, and labels. Metaformer (Diao et al., 2022) feeds a sequence of image patches and metadata to a Transformer model for their fusion. Additionally, these methods assume a uniform availability of metadata for all images, which is often not the case in reality due to data scarcity. For instance, the camera trap location coordinates may not be available in some cases due to privacy and security reasons. In our work, we do not assume such uniform availability and build the multimodal KG using available metadata.
@@
Apart from the metadata, external sources of knowledge are also used in image classification. For instance,  Jayathilaka et al. embed each class as a vector based on a hierarchy derived from WordNet (Miller, 1995). Alsallakh et al. develop a class hierarchy-aware CNN for image classification on ImageNet. Similarly, Bertinetto et al. and Zhang et al. design hierarchy-aware objectives to incorporate taxonomy in image representations. BioCLIP (Stevens et al., 2024) verbalizes the taxonomic hierarchy to train a CLIP-style foundational model for species classification across plants, animals, and fungi. Marino et al. represent images as local subgraphs of Visual Genome (Krishna et al., 2017). In contrast, COSMO constructs a global KG with both metadata and external knowledge, e.g., taxonomy information from Open Tree Taxonomy, and approaches image classification as link prediction within the KG. Our novel formulation is flexible in handling data scarcity of metadata and enables reasoning over diverse relationships present in the KG.
@@
KG Link Prediction. Most real-world KGs are incomplete. The task of link prediction or knowledge graph completion (KGC) tries to infer missing links given the observed ones. Early approaches for link prediction range from translation-based models (Bordes et al., 2013; Lin et al., 2015) and semantic matching models (Nickel et al., 2011; Yang et al., 2015) to the ones that leverage neural networks like feedforward neural networks (Dong et al., 2014), CNNs (Dettmers et al., 2018; Nguyen et al., 2018), and Transformer-based models (Yao et al., 2019; Chen et al., 2021; Saxena et al., 2022). These methods use a parameterized scoring function based on learned entity and relation embeddings to calculate the plausibility of a particular triplet. However, it could be challenging to fully encode the rich semantic information of KGs into such shallow embeddings. To mitigate this, Schlichtkrull et al., Vashishth et al., Yu et al., Pahuja et al. use graph neural networks (GNNs) to encode the rich neighborhood context of entities for link prediction. In our framework, we employ a global multimodal KG, which consists of biological taxonomy and metadata, as the context to enhance OOD generalization.
@@
Multimodal KG Reasoning. Multimodal KGs extend traditional KGs by including entities of different modalities such as categorical data, images, numerical data, etc. KBLRN (García-Durán and Niepert, 2018) is a pioneering work in multimodal KG reasoning that uses extra information in the form of relational and numerical features for multimodal KG reasoning. Similarly, IKRL (Sergieh et al., 2018) proposes a fusion of linguistic and visual information with structured information for link prediction. MKBE (Pezeshkpour et al., 2018) constructs a multimodal KG using numerical, image, and textual information, treating them as entities instead of auxiliary features, for the link prediction task. MR-GCN (Wilcke et al., 2021) further extends it by including support for more modalities, e.g., numerical, temporal, textual, visual, and spatial predicate links in the multimodal KG. To provide a more expressive way for interaction between different modalities, IMF (Li et al., 2023) uses bilinear pooling to fuse multiple modality features and trains it using contrastive learning on the contextual entity representations. Our work leverages link prediction in a multimodal KG to enable out-of-distribution generalization for species classification in camera traps.
3. Methodology
3.1. Preliminaries
@@
Problem Setup. The task entails species recognition for camera trap images amidst distribution shifts. The training and test sets comprise images obtained from disjoint camera traps, enabling the evaluation of out-of-domain (OOD) generalization. During training, we use the multimodal KG to train our model, while we use just the image to make predictions for inference. The goal is to learn visual representations robust to distribution shifts by leveraging the rich structural and semantic information provided by the multimodal knowledge graph.
3.2. Building the Multimodal KG
@@
  • Taxonomy: The taxonomy forms the core of the multimodal knowledge graph, connecting distinct species to higher-order taxa. For iWildCam2020-WILDS, we obtain the phylogenetic taxonomy corresponding to the species of interest from Open Tree Taxonomy (OTT)  (OpenTreeofLife et al., 2019) and manually link it to the species in the dataset. For the Snapshot Mountain Zebra dataset, we utilize the iNaturalist taxonomy (Horn et al., 2018) mapping provided by www.lila.science.
  • Location: The camera trap images are associated with the GPS coordinates of their source cameras. For the iWildCam2020-WILDS dataset, this metadata is available for a portion of the images (67%) and is obfuscated within 1 km. for privacy reasons. Animals demonstrate a preference for particular habitats; thus, the location context attribute is useful for species recognition.
@@
Figure 1 presents a schematic representation of various contexts in a multimodal KG. For location, time, and taxonomy attributes, the corresponding RDF triplets can be represented as (<image>, location, <GPS co-ordinate>), (<image>, time, <timestamp>), and (<taxon_1>, parent, <taxon_2>), respectively.
3.3. Model Architecture
We use DistMult (Yang et al., 2015), a strong baseline on KGE benchmarks, as our backbone KG embedding model. [^2] Recent work (Ruffinelli et al., 2020) showed that simple baselines like DistMult outperform more sophisticated neural network baselines when trained properly. Note that COSMO is a general framework that can leverage a variety of KG embedding models proposed in the literature. DistMult minimizes a bilinear scoring function between the entity embeddings of the subject and object entities. For a given triplet $(h,r,t)$, the scoring function of DistMult is defined as:
(1) $\psi(h,r,t)=\bm{h}^{T}\bm{W}_{r}\bm{t}=\sum_{i=1}^{d}\bm{h}_{i}\cdot diag(\bm{W}_{r})_{i}\cdot\bm{t}_{i}$